### v2

# Contest for system observability as an infinitely repeated game

Abstract
This paper studies a system security problem in the context of observability based on a two-person non-cooperative infinitely repeated game. Both the attacker and the defender have means to modify the dimension of the unobservable subspace, which is set as the value function. Utilizing tools from geometric control, we construct the best response sets considering one-step and two-step optimality respectively to maximize or minimize the value function. We find that the best response sets are not single-valued maps, resulting in a variety of game outcomes. In addition, we establish a necessary and sufficient condition for Nash equilibrium. The game analysis reveals that the defender plays a pivotal role in steering the game towards equilibrium, while the attacker can disrupt the equilibrium by sacrificing immediate gains.
keywords:
Observability, Linear system, Repeated games, Nash equilibrium, Geometric control
thanks: This work was supported in part by the National Natural Science Foundation of China with No. 62373245, in part by the National Key Research and Development Program of China with No. 2023YFB4706800, and in part by the Program of Shanghai Education Commission, China. Partial results of this paper were presented at IFAC World Congress 2023. Corresponding author Lin Wang.
1 Introduction
In recent years, more and more attention has been paid to the security of control systems. Remote sensors are vulnerable to attacks, which intend to deteriorate system performance by manipulating the transmitted data while remaining stealthiness [17]. Those attacks may have severe consequences. For example, in the cyber attack on the Ukrainian power grid in 2015, the attacker operated several of the circuit breakers in the grid and jammed the communication network to cause a large scale blackout and keep the operators unaware [21]. If attack can be found in time, the damage can be reduced.
In this paper, we want to study attacks against observability. Observability is a critical aspect of system performance. When observability is destroyed, not only the aforementioned attacks are harder to be detected, but there are more severe consequences. For example, observers relying on observability of the system become unusable [4] and operators are unable to accurately recover the true state of the system [20]. Plus, systems that lose observability may be more susceptible to various forms of stealthy attacks [11]. Researchers have recognized the significance of study on observability under attack in control systems. Regarding state reconstruction after attacks, observability of linear systems has been studied for the case where an attacker modified some outputs [6] (along with similar findings by [7] and [19]). If more than half of the sensors were attacked, accurately reconstructing the initial state would become impossible. In [14], the concept of eigenvalue observability is introduced to estimate locally undetectable states caused by attacks when a single node can exchange information with its neighbours. Later the same authors develop a fully distributed algorithm that successfully reconstructed the system state despite the presence of sensor attacks within the network [15]. In [25], the robustness of observability of a linear time-invariant system under sensor failures is studied from a computational perspective.
However, it should be emphasized that the above research focuses on whether the system can reconstruct the initial state qualitatively. They did not study dimension change of the unobservable subspace quantitatively, which can greatly impact the effectiveness of stealthy attacks. In this context increasing the dimension of the unobservable subspace can facilitate covert attacks, since the attack vector space for stealthiness is getting bigger [9], [26], [13]. In [9] and [26], subspace methods are used to construct undetectable attacks, yet without altering the dimension of the unobservable subspace. In [13], the attacker enhances the effect of stealthy attack by masking sensor signals to increase the dimension of the unobservable subspace.
In recent years, the game approach has shed new light on the system security problem [28], [10]. When the attacker and the defender have limited information, the game becomes partially observable. In [8], a partially observable stochastic game is studied and the authors give a representation of uncertainty encountered by the defender. In [27], an $\epsilon$-Stackelberg partially observable game model is built to prevent state information leakage. However, observability is considered as a game setting in the above research and few game research directly involves confrontation on the observability between attackers and defenders. In [12], a game approach is used to study the stealthy attack problem in which the attacker masked the sensors and the defender reinforced the sensors. However, in the utility of the players, observability is abstracted into different values without considering the relationship between observability and structural matrices of the system [12].
Repeated game is a game model where participants engage in the same basic game for many times. In the system security field, repeated game is a potent tool to analyze the dynamic interplay between attackers and defenders and help both sides to design strategies in their favor. A finite repeated security game is constructed in [16] to study how an attacker manipulates attack data to mislead the defender, thereby influencing the defender’s learning process to make the game more favorable to itself. In [2], an infinitely repeated game and cooperation method are designed to detect malicious nodes and improve energy efficiency. In response to the non-convex infinitely repeated game problem, algorithms are constructed to select both the optimal security strategies that necessitate monitoring and those strategies that do not [1].
Additionally, most existing game models use continuous value functions such as quadratic functions, and equilibrium solutions are obtained through methods such as dynamic programming in [23] and Q-learning in [18]. However, these methods are not suitable for discrete value functions, such as dimension of the unobservable subspace.
To address these limitations, an infinitely repeated game between the attacker and the defender is proposed, where the attacker aims to destroy system observability by maximizing the dimension of the unobservable subspace, while the defender tries to protect it. The main contributions are as follows.
(1) By taking the dimension of unobservable subspace as the value function, derivations and algorithms to maximize or minimize the dimension are given. The game antagonism based on observability is quantitatively studied.
(2) We investigate the impact of considering one-step and two-step optimality on game outcomes. We reveal that the best response sets are multi-valued, leading to diverse outcomes. Additionally, we give the condition for Nash equilibrium applicable to both scenarios.
(3) We analyze the disparities in the upper limits of capabilities between the attacker and defender. The defender dominates the game to achieve Nash equilibrium, while the attacker can forgo short-term benefits to prevent the establishment of the equilibrium, in order to gain long-term benefits.
Based on [24], this paper further studies the infinitely repeated game by incorporating two-step optimization in addition to one-step optimization. We find different action choices in the best response sets will lead to different game results. We analyze the Nash equilibrium condition and find the different effects of the attacker and defender on game equilibrium respectively.
The rest of paper is organized as follows. In Section 2, we formulate the system model and the game problem. In Section 3, derivations and algorithms are given to get the best response sets for both players. Section 4 analyzes the equilibrium and results for the game considering one-step optimality and two-step optimality respectively. In Section 5, examples are given to illustrate various game results. Section 6 is a brief conclusion. A summary of notations is provided in Table 1.
Table 1: Notations
Notations Definitions
$\mathbb{R}$ set of real numbers
$\mathbb{R}^{n}$ set of n-dimensional real vectors
$\mathbb{R}^{m\times n}$ set of $m\times n$-dimensional real matrices
$\operatorname{Im}A,A\in\mathbb{R}^{m\times n}$ image space, $\{v\in\mathbb{R}^{m}:v=Aq,\forall\,q\in\mathbb{R}^{n}\}$
$\operatorname{Ker}A,A\in\mathbb{R}^{m\times n}$ kernel space, $\{w\in\mathbb{R}^{n}:Aw=0\}$
$\operatorname{pinv}(A)$ pseudo inverse, $\left(A^{\top}A\right)^{-1}A^{\top}$
$\operatorname{Col}_{k}(A)$ the $k$-th column of matrix $A$
$\delta_{n}^{k}$ $\operatorname{Col_{k}}\left(I_{n}\right)$, the $k$-th column of $I_{n}$
$\oplus$ direct sum
2 Problem formulation
2.1 Modeling of the system
Consider the following linear system
$\begin{split}\dot{x}&=Ax+B_{d}u_{d}+B_{a}u_{a},\\ y&=Cx,\end{split}$ (1)
where $x\in\mathbb{R}^{n_{0}}$, and $y\in\mathbb{R}^{m}$ are state and output of the system respectively; $A\in\mathbb{R}^{n_{0}\times n_{0}}$ is system matrix and $C\in\mathbb{R}^{m\times n_{0}}$ is output matrix. $u_{d}\in\mathbb{R}^{m}$ and $u_{a}\in\mathbb{R}^{k}$ are two input channels controlled by the defender and the attacker respectively.
In this paper, it is assumed that the attacker wants to destroy observability and maximize the unobservable subspace by injecting feedback-data using $u_{a}$. To the contrary, the system defender wants to protect the system observability and minimize the unobservable subspace via the input $u_{d}$. Let us first recall $\mathcal{V}^{*}$ space and friend matrices.
Definition 1.
$\mathcal{V}$ is a controlled invariant (or $(A,B)$-invariant) subspace if there exists a matrix $F$ such that $(A+BF)\mathcal{V}\subseteq\mathcal{V}$. Such an $F$ is called a friend matrix of $\mathcal{V}$ and we denote the set of friend matrices by $\mathcal{F}(\mathcal{V})$. Define controlled invariant subspaces contained in space $Z$ as $S(Z)$. In $S(\operatorname{Ker}\,C)$, there is a maximal one which is denoted as $\mathcal{V}^{*}$.
In this paper, we assume the square system $(A,B_{d},C)$ has the relative degree $(r_{1},...,r_{m})$, which reflects the order of differentiation needed in order to have the input $u_{d}$ explicitly appearing in the output $y$. It is also assumed that the attacker is employing a specific kind of stealthy attack, namely a zero-dynamics attack [22], which requires $\mbox{Im}\,B_{a}\subseteq\mathcal{V}^{*}$, where $\mathcal{V}^{*}$ is the maximal $(A,B_{d})$-invariant subspace in Ker $C$.
With the relative degree, it is well known that after a coordinate change $x\rightarrow\left(\begin{array}[]{l}z\\ \xi\end{array}\right)$, where $\xi=\left[\xi_{1}^{1},\xi_{2}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\xi_{2}^{m},\cdots,\xi_{r_{m}}^{m}\right]^{\top}$, system (1) can be rewritten as the normal form
$\dot{z}$ $=Nz+E\xi+B_{a}^{\prime}u_{a},$ (2)
$\dot{\xi}_{1}^{i}$ $=\xi_{2}^{i},$
[…]
$\dot{\xi}_{r_{i}-1}^{i}$ $=\xi_{r_{i}}^{i},$
$\dot{\xi}_{r_{i}}^{i}$ $=R_{i}z+S_{i}\xi+c_{i}A^{r_{i}-1}B_{d}u_{d},$
$y_{i}$ $=\xi_{1}^{i},$ (4)
where $i=1,\ldots,m,y=[y_{1},y_{2},\cdots,y_{m}]^{\top}$, $B_{a}^{\prime}$ is determined by $B_{a}$. Since the attacker can in essence only change the zero dynamics of the system, the attack is a zero-dynamics attack.
The complete evolution equation of $\xi_{r_{i}}^{i}$ can be written as
$[\dot{\xi}_{r_{1}}^{1^{\top}},\dot{\xi}_{r_{2}}^{2^{\top}},\cdots,\dot{\xi}_{r_{m-1}}^{m-1^{\top}},\dot{\xi}_{r_{m}}^{m^{\top}}]^{\top}=Rz+S\xi+Lu_{d},$ (5)
where $R\in\mathbb{R}^{m\times(n_{0}-s)}$, $S\in\mathbb{R}^{m\times s}$, $L\in\mathbb{R}^{m\times m}$, $s=\sum_{i=1}^{m}r_{i}$. We define the feedback controls as
$u_{d}=K_{d}\xi+U_{d}z+u_{0},\,u_{a}=K_{a}\xi+U_{a}z,$ (6)
where $K_{d}\in\mathbb{R}^{m\times s}$, $U_{d}\in\mathbb{R}^{m\times(n_{0}-s)}$ are determined by the defender; $K_{a}\in\mathbb{R}^{k\times s}$, $U_{a}\in\mathbb{R}^{k\times(n_{0}-s)}$ are determined by the attacker; $u_{0}$ is the input that maintains the normal operation of the system and is controlled by a basic controller, which is illustrated in Fig. 1(a). Then
$[\dot{\xi}_{r_{1}}^{1^{\top}},\cdots,\dot{\xi}_{r_{m}}^{m^{\top}}]^{\top}=(R+LU_{d})z+(S+LK_{d})\xi+Lu_{0}.$ (7)
(a) Model in the normal form
(b) Model in the observability equivalent system
Figure 1: Comparison of the model in the normal form and in the observability equivalent system.
Considering the scenario where $u_{0}$ does not incorporate state feedback, it consequently does not affect the observability of the system. Thus $u_{0}$ is omitted in the following derivation of the observability equivalent system. Define
$R+LU_{d}\triangleq{H}.$
Since $L$ is a non-singular matrix according to the definition of relative degree, $\forall{H}\in\mathbb{R}^{m\times n}$, there exists $U_{d}=L^{-1}\left({H}-R\right)$. Thus the defender can completely control ${H}$. If the system is observable, when $y\equiv\mathbf{0}$, there is $\left(\begin{array}[]{l}z\\ \xi\end{array}\right)\equiv\mathbf{0}$. Consider $y\equiv\mathbf{0}$, according to the coupling relationship between $y_{i}$ and $(\xi_{1}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\cdots,\xi_{r_{m}}^{m})=\xi^{\top}$, we have $\xi\equiv\mathbf{0}$. Thus we only need to prove $z\equiv\mathbf{0}$, where $z$ satisfies
$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \mathbf{0}&=Hz.\end{split}$ (8)
Define $Hz=\hat{y}$ and $n_{0}-s=n$. The condition for completely observable becomes: if $\hat{y}\equiv\mathbf{0}$, there is $z\equiv\mathbf{0}$, which is also the condition for the following system to be observable:
$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \hat{y}&=Hz,\end{split}$ (9)
where $z\in\mathbb{R}^{n}$ and $\hat{y}\in\mathbb{R}^{m}$ are state and output of the observability equivalent system respectively; $N\in\mathbb{R}^{n\times n}$ is system matrix and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ is input matrix. Now the problem becomes that the attacker wants to damage system observability using state feedback control $U_{a}$, while the defender wants to protect system observability by controlling $H$, which is illustrated in Fig. 1(b).
2.2 Game formulation
Define an infinitely repeated game as a tuple $(N,A,J)$. $N=\{attacker(a),defender(d)\}$ is the set of players. $A=A^{a}\oplus A^{d}$ is the action set of players, where $A^{a}=\{U_{a}\in\mathbb{R}^{k\times n}\},A^{d}=\{H\in\mathbb{R}^{m\times n}\}$. $J=\{J^{a},J^{d}\}$ is the utility function of the attacker and the defender. Define the utility function of the attacker in epoch $i$ as
$J^{a}(U_{ai},H_{i})$ $=\operatorname{dim\,Ker}\!\!\left[\begin{array}[]{c}H_{i}\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)\\ \vdots\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)^{n-1}\end{array}\right]$
$\overset{\triangle}{=}\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}),$
where $\Omega$ is the observability matrix of system (9), $U_{ai}$ and $H_{i}$ are actions of players in epoch $i$. As for the defender,
$J^{d}(U_{ai},H_{i})$ $=-J^{a}(U_{ai},H_{i}).$ (10)
Define the value function $\Phi$ in epoch $i$ as the dimension of unobservable subspace
$\Phi(U_{ai},H_{i})=\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}).$ (11)
Thus the attacker aims to maximize the value function, while the defender seeks to minimize. Two players update actions asynchronously and an epoch is defined once a player acts. Repetitive games can effectively reflect the interaction between the attacker and defender, guiding both sides in designing strategies to gain an advantage in this tug-of-war relationship.
3 The best response sets
In this section, we will give derivations and algorithms to get the best response sets of the attacker and defender respectively. We assume that matrices $N,B_{a}^{\prime}$ and players’s actions $U_{a},H$ are public knowledge.
3.1 The best response set of the attacker
The attacker aims to maximize the value function by controlling $U_{a}$, i.e.,
$U_{a}^{*}=\!arg\!\max_{U_{a}\in\mathbb{R}^{k\times n}}\!\operatorname{dim\,Ker}\Omega(U_{a},H).$ (12)
Denote $\mathcal{V}\triangleq\operatorname{Ker}\,\Omega$. It is apparent that $\mathcal{V}$ is a controlled invariant subspace contained in $\operatorname{Ker}H$. Among all controlled invariant subspaces contained in $\operatorname{Ker}H$, there is a maximal one denoted as $\mathcal{V}^{*}$, thus,
$\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}.$ (13)
A lemma is given to find $\mathcal{V}^{*}$.
Lemma 1.
[3] Let  $\mathcal{V}_{0}=\operatorname{Ker}H$ and define, for $\ i=0,1,2,$,
$\ \mathcal{V}_{i+1}=\left\{x\in\operatorname{Ker}H\mid Nx\in\mathcal{V}_{i}+\operatorname{Im}B_{a}^{\prime}\right\}.$ (14)
Then $\mathcal{V}_{i+1}\subset\mathcal{V}_{i}$. There exists $q\in\mathbb{R},\,q\leq\operatorname{dim\,}\mathcal{V}_{0}$, $\mathcal{V}_{q+1}=\mathcal{V}_{q}=\mathcal{V}^{*}$.
By classical results on maximal controlled invariant subspace, we can conclude that $U_{a}\in\mathbb{R}^{k\times n}$ maximizes the dimension of unobservable subspace if and only if it is a friend matrix of $\mathcal{V}^{*}$, i.e.,
$U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}),$ (15)
where $\mathcal{F}(\mathcal{V}^{*})$ is defined in Definition 1. In Algorithm 1, a way using pseudo inverse to find a friend matrix is given. For most cases, the friend matrix is not unique. Since in our case, all friend matrices are equivalent, so we use the pseudo inverse to get the explicit form for $U_{a}^{*}$. The calculation steps of $U_{a}^{*}$ is summarized in Algorithm 1.
Algorithm 1 Algorithm for the attacker to maximize the unobservable subspace
  Set $\operatorname{Im}V_{0}=\operatorname{Ker}H$. Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{0}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$ Calculate $\operatorname{Im}V_{1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
  while $\operatorname{Im}V_{i}\neq\operatorname{Im}V_{i+1}$ do
     Set $V_{i}=V_{i+1}$. Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{i}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$. Calculate $\operatorname{Im}V_{i+1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
  end whileSet $V=V_{i}=\left[\begin{array}[]{llll}v_{1}&v_{2}&\ldots&v_{r}\end{array}\right]$.Calculate $[X;U]=\text{pinv}\left(V\quad B_{a}^{\prime}\right)NV$, choose the last $\quad(n-r)$ columns as $U$.Calculate $U_{a}=-U\text{pinv}(V)$.
3.2 The best response set of the defender
The defender aims to minimize the value function by controlling $H$, i.e.,
${H^{*}}=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),$ (16)
which is equal to $\arg\max_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Im}\Omega^{T}$,where $\Omega^{T}\!=[H^{\top}\!,\!\left(N+B_{a}^{\prime}U_{a}\right)^{\top}\!H^{\top},\!\cdots,\left(N+B_{a}^{\prime}U_{a}\right)^{(n-1)\top}\!H^{\top}]$ can be viewed as controllability matrix of the dual system of (9):
$\dot{\bar{z}}=\bar{A}\bar{z}+\bar{B}\bar{u},$ (17)
where $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top},\bar{B}=H^{\top}$. Thus the best response set (16) becomes
$\bar{B}^{*}=\arg\max_{\bar{B}}\operatorname{dim\,Im}\left[\begin{array}[]{lll}\bar{B},\!\!&\bar{A}\bar{B},\cdots,\bar{A}^{n-1}\bar{B}\end{array}\!\!\right].$ (18)
The problem becomes how to choose $\bar{B}$ to make the dual system controllable. Find similar transformation matrix $T\in\mathbb{R}^{n\times n}$ which makes $\bar{A}$ become Jordan normal form $J$, i.e., $J=T^{-1}\bar{A}T$. And $\bar{B}$ becomes $\hat{B}=T^{-1}\bar{B}$. For the above $J$, let its $l$ eigenvalues be: $\lambda_{1}$ (algebraic multiplicity: $\sigma_{1}$, geometric multiplicity: $\alpha_{1}$), $\lambda_{2}$ (algebraic multiplicity: $\sigma_{2}$, geometric multiplicity: $\alpha_{2}$), $\cdots$, $\lambda_{l}$ (algebraic multiplicity: $\sigma_{l}$, geometric multiplicity: $\alpha_{l}$). Assume $\lambda_{i}\neq\lambda_{j},\forall i\neq j$ and $\sigma_{1}+\sigma_{2}+\cdots+\sigma_{l}=n$.
Thus we have
$J=J\left(\lambda_{1}\right)\oplus J\left(\lambda_{2}\right)\oplus\cdots\oplus J\left(\lambda_{l}\right),$ (19)
$\hat{B}=\left[\hat{B}_{1}^{\top},\hat{B}_{2}^{\top},\cdots,\hat{B}_{l}^{\top}\right]^{\top},$ (20)
where $\oplus$ is the direct sum of matrices,
$J(\lambda_{i})=\begin{bmatrix}{J}_{1}(\lambda_{i})&&&\\ &{J}_{2}(\lambda_{i})&&\\ &&\ddots&\\ &&&{J}_{\alpha_{i}}(\lambda_{i})\end{bmatrix},\hat{B}_{i}=\begin{bmatrix}\hat{{B}}_{i1}\\ \hat{{B}}_{i2}\\ \vdots\\ \hat{{B}}_{i\alpha_{i}}\end{bmatrix},$
for $i=1,2,\cdots,l$, where
$J_{k}(\lambda_{i})=\begin{bmatrix}\lambda_{i}&1&&&\\ &\lambda_{i}&1&&\\ &&\ddots&\ddots&\\ &&&\ddots&1\\ &&&&\lambda_{i}\end{bmatrix},\hat{{B}}_{ik}=\begin{bmatrix}\hat{{b}}^{ik}_{1}\\ \hat{{b}}^{ik}_{2}\\ \vdots\\ \hat{{b}}^{ik}_{r_{ik}-1}\\ \hat{{b}}^{ik}_{r_{ik}}\end{bmatrix},$
for $k=1,2,\cdots,\alpha_{i}$, where $J_{k}(\lambda_{i})\in\mathbb{R}^{r_{ik}\times r_{ik}},\hat{{B}}_{ik}\in\mathbb{R}^{r_{ik}\times m}$. $\sum_{k=1}^{\alpha_{i}}r_{ik}=\sigma_{i}$. Recall the following Lemma.
Lemma 2.
[5] For the Jordan normal form of linear system, the necessary and sufficient condition for complete controllability is
$\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\forall i=1,2,\cdots,l,$ (21)
which means the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are linearly independent.
However, if there is a geometric multiplicity of an eigenvalue that exceeds the number of columns of $\hat{B}$, i.e., $\exists\alpha_{i}>m$, it is impossible for the system to be fully controllable. In this case, we give a condition to maximize the controllable subspace.
Proposition 3.1.
$\hat{B}\in\mathbb{R}^{n\times m}$ maximizes the dimension of the controllable subspace if and only if the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
$\left\{\begin{array}[]{cc}\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]=m,\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (22)
where $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$.
Proof 3.2.
For $m\geq\alpha_{i}$, (sufficiency and necessity) according to Lemma 2, it is the sufficient and necessary condition for complete controllability. Thus the dimension of controllable subspace is maximized.
For $m<\alpha_{i}$, (sufficiency) if $\operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]$ $=m$, the dimension of controllable subspace $\operatorname{dim\,Im}\Omega^{\top}=\sum_{p=1}^{m}r_{ij_{p}}.$ Because $\{j_{1},j_{2},\cdots,j_{m}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$, the dimension of controllable subspace is maximized. (Necessity) Given the symmetry in the argument, the necessity part is omitted.
With $\hat{B}$ which satisfies Proposition 3.1, the best response for the defender is
[…]
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Based on Proposition 3.1, we can obtain the lower bound of the value function minimized by the defender.
Corollary 3.3.
For ${H}\in\mathbb{R}^{m\times n}$, define the set $\mathcal{I}$ which includes subscripts for the eigenvalues of $(N+B_{a}^{\prime}U_{a})$ whose geometric multiplicity is larger than $m$, i.e., =$\mathcal{I}=\{i\mid\alpha_{i}>m\}$. Then $\forall U_{a}\in\mathbb{R}^{k\times n}$,
$&\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\!\operatorname{dim\,Ker}\Omega(U_{a},H)\!=\!n\!-\!\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}}),\,\text{for }\mathcal{I}\neq\emptyset,\\ &\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{a},H)=0,\,\text{for }\mathcal{I}=\emptyset,\end{aligned}\right.$ (24)
where $\{r_{qj_{1}},r_{qj_{2}},\cdots,r_{qj_{m}}\}$ are the dimensions of $m$-th largest Jordan blocks for eigenvalue $\lambda_{q}$.
Based on Proposition 3.1, we can also give a specific way to construct $\hat{{B}}$.
Corollary 3.4.
If the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
$\left[\!\begin{array}[]{c}\hat{b}_{r_{i1}}^{i1}\\ \hat{b}_{r_{i2}}^{i2}\\ \vdots\\ \hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}}\end{array}\!\right]\!\!=\!\!\left\{\begin{array}[]{cc}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right],\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right],\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (25)
where  $\delta_{\alpha_{i}}^{j_{p}}=\operatorname{Col}_{j_{p}}\left(I_{\alpha_{i}}\right)$, $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$ and other rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are zero rows, then the controllable subspace is maximized.
The calculation steps of $H^{*}$ based on Corollary 3.4 is summarized in Algorithm 2.
Algorithm 2 Algorithm for the defender to minimize the unobservable subspace
  Set $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top}.$ Calculate Jordan normal form of $\bar{A}:T^{-1}\bar{A}T=J$. Calculate geometric multiplicity $\alpha_{i}$ of $J$’s eigenvalue $\lambda_{i},i=1,2,\cdots,l$ ; Calculate dimensions $\{r_{i1},r_{i2},\cdots,r_{i\alpha_{i}}\}$ of Jordan blocks of $\lambda_{i},i=1,2,\cdots,l$; Calculate $\{j_{1},j_{2},\cdots,j_{m}\}$ which are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$. Set $\hat{B}\in\mathbb{R}^{n\times m}$ to be an all-zeros matrix. Update $\hat{B}$ according to (25). Calculate $H=\hat{B}^{\top}T^{\top}$.
4 Equilibrium analysis
Based on the above best response sets, we further analyze the equilibrium of the game considering one-step optimality and two-step optimality respectively in the subsections 4.1 and 4.2. In the subsection 4.3, we give equilibrium analysis which is suitable for both cases.
4.1 Game based on one-step optimality
The infinitely repeated game considering one-step optimality means that players only consider the best of the current epoch. Define the best response sets for the attacker and the defender in this case as $BR1^{a}$ and $BR1^{d}$, i.e.,
$BR1^{a}=\arg\max_{{U_{a}}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),$ (26)
$BR1^{d}=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (27)
Fig. 2 shows the sequence of actions. The solutions of those two best response sets have been discussed in subsections 3.1 and 3.2, i.e.,
$BR1^{a}=\mathcal{F}(\mathcal{V}^{*}),$ (28)
where $\mathcal{F}(\mathcal{V}^{*})$ is the friend matrices of $\mathcal{V}^{*}$, and $\mathcal{V}^{*}$ is the maximal controlled invariant subspace in $\operatorname{Ker}H$. And
[…]
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Both sets are not single-valued maps and Algorithm 1 (or Algorithm 2) only determines a special $U_{a}$ (or $H$) in $BR1^{a}$ (or $BR1^{d}$). If a player chooses a different action in the best response set, the game result will be different.
Figure 2: Sequence of actions considering one-step optimality.
Define the Nash equilibrium.
Definition 2.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be the Nash equilibrium of the one-step optimal game, if $U_{a}^{*}\in BR1^{a},H^{*}\in BR1^{d}$.
Then we give a necessary and sufficient condition for the Nash equilibrium.
Theorem 4.5.
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium of the one-step optimal game if and only if $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)\!=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})\!=\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$.
Proof 4.6.
(Sufficiency) Because $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $(U_{a}^{*},H)=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$, $H^{*}\in BR1^{d}$. Since $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*}),U_{a}^{*}$ is a friend matrix of $\mathcal{V}^{*}(H^{*})$ and $U_{a}^{*}\in BR1^{a}$. Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium. (Necessity) Given the symmetry in the argument, the necessity part is omitted.
Remark 4.7.
Theorem 4.5 is effective in determining when the Nash equilibrium is reached in dynamic game process, as it only requires verifying if $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)$ $=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})$, i.e., whether the value function after the defender chooses $H^{*}$ is equal to $\operatorname{dim}\mathcal{V}^{*}(H^{*})$. If this condition is satisfied, since $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$ always holds when both players consider one-step optimality, the condition in Theorem 4.5 is satisfied and the game reaches Nash equilibrium.
Figure 3: An example of the evolution of best response sets and actions. Coordinate axis represents game epochs. Different shapes above the axis symbolize different best response sets per epoch. Differently colored points represent different actions; those below the axis represent the actions chosen in each epoch.
Next we want to analyze the game result in dynamic game process. Since the best response sets for both players are not single-valued maps, we need to add assumptions on the choices for both players:
Assumption 1: The attacker (or the defender) prefers to keep the action unchanged if the last action also belongs to the best response set in this epoch.
Assumption 2: Without violating Assumption 1, if the best response set is the same in different game epochs, the attacker (or defender) consistently chooses the same action as the first time.
To further illustrate the above assumptions, Fig. 3 represents an example of the evolution of the best response sets and actions for one player, who updates its action every two epochs. In epoch 1, the player chooses an action (the red point) in the best response set (the square pattern). Then in epoch $k+2$, the player has the same best response set as that in epoch 1 and chooses the same action as that in epoch 1 according to Assumption 2. Afterwards, in epoch $m+2$, though the player has the same best response set as that in epoch 1, it keeps the action unchanged (the blue point), because the last action (the blue point) also belongs to the best response set in this epoch and Assumption 1 is prioritized over Assumption 2.
Assumptions 1 and 2 frequently manifest in real-world applications. Assumption 1 can save energy used in changing actions, while Assumption 2 fixes the rule to select a solution in a multi-valued map (such as choosing the one with the smallest norm length).
Based on these assumptions, we can easily get the following three corollaries which represent different game results.
Corollary 4.8.
For infinitely repeated game considering one-step optimality, the value function is locked, i.e., $\exists l\in\mathbb{N}$ and $\gamma\in\mathbb{R}$, in epoch $i\geq l,\operatorname{dim\,Ker}\Omega(U_{ai},H_{i})=\gamma$, if and only if players’ strategy profile in epoch $l$ $(U_{al},H_{l})$ is a Nash equilibrium.
Remark 4.9.
This corollary is nontrivial when the best response sets are not single-valued maps, because different choices in the best response set affect the upper bound of the subsequent action of another player. Therefore we need the assumptions 1-2.
Corollary 4.10.
If in epoch $l$, players’ strategy profile $(U_{al},H_{l})$ is not a Nash equilibrium, the value function oscillates and the amplitude equals to $\operatorname{dim\,}\mathcal{V}^{*}(H_{l})-\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{al},H).$
Corollary 4.11.
If the attacker or the defender chooses the same action in epoch $i$ and epoch $j,i<j$ with $(j-i)$ even, then the value function and players’ actions will evolve in a circle and $(j-i)$ is a multiple of the number of epochs in a minimal circle.
Proof 4.12.
Assume $U_{ai}=U_{aj}$. In epoch $j$, $(N+B_{a}^{\prime}U_{aj})$ returns to the same value as $(N+B_{a}^{\prime}U_{ai})$. Thus in epoch $(j+1)$, the $BR1^{d}$ is the same as $BR1^{d}$ in epoch $(i+1)$. According to Assumption 2, the defender will choose the same action in epoch $(j+1)$ as in epoch $(i+1)$. (Here we do not need to consider the scenario where Assumption 1 is satisfied, because in this case, the game reaches Nash equilibrium, which is also a circle.) Similarly, at each subsequent epoch, the attacker (or defender) will take the same action as $(j-i)$ epochs before, which leads to the circle of the value function and players’ actions. Thus $(j-i)$ is a multiple of the number of epochs in a minimal circle. If $H_{i}=H_{j}$, the proof is similar.
4.2 Game based on two-step optimality
The best response sets considering one-step optimality are not single-valued maps. Different initial preferences for choices based on Assumption 2 can lead to different game outcomes, which is showed in Section 5 Illustrative examples. This is because on the one hand, different values of $H$ in the set $BR1^{d}$ have different values of $\operatorname{dim\,}\mathcal{V}^{*}$, which determines the upper bound of $U_{a}$ to maximize the value function in the next epoch. On the other hand, different values of $U_{a}$ in the set $BR1^{a}$ will influence the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$, which influences the lower bound of $\min_{H}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ in the next epoch. Thus $\operatorname{dim\,}\mathcal{V}^{*}$ and the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$ are taken into consideration in this section.
Formulate the problem based on two-step optimality, which means two players are smart enough to choose an action in $BR1^{a}$ or $BR1^{d}$ that gives the opponent the worst outcome in the next step. There are two steps to find the best response sets following two-step optimality: $(i)$ find $BR1^{a}$ or $BR1^{d}$; $(ii)$ in $BR1^{a}$ or $BR1^{d}$, find the set of actions that make the opponent get the worst outcome in the next epoch.
The best response set for the attacker built on two-step optimality is
$\begin{split}BR2^{a}&=\arg\max_{U_{a}\in BR1^{a}}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ &=\arg\max_{U_{a}\in BR1^{a}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].\end{split}$ (30)
where $\text{MGM}(N+B_{a}^{\prime}U_{a})$ is the maximum geometric multiplicity associated with the eigenvalues of $(N+B_{a}^{\prime}U_{a})$. This result can be directly derived from Corollary 3.3. Define $BR2X^{a}$ as
$BR2X^{a}=\arg\max_{U_{a}\in\mathbb{R}^{k\times n}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].$ (31)
$BR2X^{a}$ is not the best response set for the attacker considering two-step optimality, since it does not guarantee that the value function of the current epoch is maximized.
Because $\max_{U_{a}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}(H)$, the best response set for the defender following two-step optimality is
$\begin{split}BR2^{d}=&arg\min_{H\in BR1^{d}}\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ =&arg\min_{H\in BR1^{d}}\operatorname{dim\,}\mathcal{V}^{*}(H).\end{split}$ (32)
Define set $BR2X^{d}$ as
$BR2X^{d}=\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H).$ (33)
$BR2X^{d}$ does not ensure the value function of the current epoch is minimized; hence, it is not the best response set for the defender considering two-step optimality.
Figure 4: Sequence of actions following two-step optimality.
Remark 4.13.
Actually, $BR2X^{d}$ and $BR2X^{a}$ are Stackelberg game equilibrium in two different scenarios. In a Stackelberg game, the leader possesses knowledge of the follower’s utility function and takes the initial action, taking into account how the follower is likely to respond. The follower, on the other hand, has the ability to observe the leader’s action and change its own action accordingly. In scenario 1, the defender is the leader and the attacker is the follower. The best response of the follower is $BR1^{a}$ and the Stackelberg equilibrium action of the leader is $BR2X^{d}$. In scenario 2, the attacker is the leader and the defender is the follower. The best response of the follower is $BR1^{d}$ and the Stackelberg equilibrium action of the leader is $BR2X^{a}$.
Fig. 4 shows the sequence of actions. Then we define the Nash equilibrium when two players consider two-step optimality.
Definition 3.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be a Nash equilibrium of the two-step optimal game if $U_{a}^{*}\in BR2^{a},H^{*}\in BR2^{d}$.
In order to get the condition for Nash equilibrium, we give the following lemma.
Lemma 3.
For the same $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)$.
Proof 4.14.
$\forall\,H_{0}\in\mathbb{R}^{m\times n}$, it is apparent that the maximal controlled invariant subspace in Ker $H_{0}$ is larger than $(N+B_{a}^{\prime}U_{a})$-invariant subspace in Ker $H_{0}$, i.e., $\operatorname{dim\,}\mathcal{V}^{*}(H_{0})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{0}).$ Thus $\forall H_{1}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{1})$. And $\forall H_{2}\in\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a},H)$, there is $\operatorname{dim\,Ker}\Omega(U_{a},H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$. Thus $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\!\geq\!\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$.
Based on this lemma, it is found that the Nash equilibrium condition in Theorem 4.5 is also valid when two players consider two-step optimality. Although the equilibrium conditions for the player considering one-step or two-step optimality are the same, the mathematical reasoning is different.
Theorem 4.15.
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium when two players consider two-step optimality if and only if, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$.
Proof 4.16.
(Sufficiency) According to Theorem 4.5, $H^{*}\in BR1^{d}$ and $U_{a}^{*}\in BR1^{a}$. We need to further prove $H^{*}\in BR2^{d}$ and $U_{a}^{*}\in BR2^{a}$. $\forall H\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma,$ where $\gamma$ is a constant for different $H\in BR1^{d}$. Because $\exists H^{*}\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})=\gamma$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ reaches the lower bound. Thus $H^{*}\in BR2^{d}$. Define $F_{1}(U_{a})=\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega(U_{a},H),$ then the upper bound of $F_{1}(U_{a})$ is $\overline{F_{1}}\!=\!\max_{U_{a}\in\mathbb{R}^{k\times n}}\!\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}$ $\Omega(U_{a},H).$ According to Lemma 3, $\overline{F_{1}}\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$. Since $F_{1}(U_{a}^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$, there must be $F_{1}(U_{a}^{*})=\overline{F_{1}}=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Because $U_{a}^{*}$ lets $F_{1}$ reaches the upper bound, $U_{a}^{*}\in BR2^{a}$. Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium under the game framework of considering two-step optimality.
(Necessity) Since $BR2^{d}\in BR1^{d}$ and $BR2^{a}\in BR1^{a}$, the rest proof is the same as Theorem 4.5.
Although the best response sets for two-step optimality are much smaller than that for one-step optimality, they still are not a singleton. Therefore, we extend Assumption 1 and 2 to the game framework considering two-step optimality. Then we can easily find that corollaries 4.8-4.11 also hold in game following two-step optimality.
4.3 General equilibrium analysis
This subsection gives conclusions that are consistent with both one-step and two-step optimality game. Firstly, we introduce a special equilibrium condition.
Theorem 4.17.
For $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, if $\exists$ $H^{*}\in\mathbb{R}^{m\times n}$ satisfies $(1)$ or $(2)$:
[…]
then $\operatorname{dim}$ $\mathcal{V}^{*}(H^{*})=0.$ And $\forall U_{a}\in\mathbb{R}^{k\times n}$ the strategy profile $(U_{a},H^{*})$ is an equilibrium.
Proof 4.18.
$(1)\text{ When }(n-m)\geq k.\text{ Since }\operatorname{Im}B_{a}^{\prime}\subseteq\operatorname{Ker}H^{*},H^{*}B_{a}^{\prime}=0.\text{ Then }\forall v\neq 0\in\operatorname{Ker}H^{*},\forall U_{a},H^{*}(N+B_{a}^{\prime}U_{a})v=H^{*}Nv\neq 0.$ Thus $(N+B_{a}^{\prime}U_{a})v\notin\operatorname{Ker}H^{*}$, which means controlled invariant subspace contained in $\operatorname{Ker}H^{*}$ is 0. Thus dim $\mathcal{V}^{*}(H^{*})=0.$
[…]
When dim $\mathcal{V}^{*}(H^{*})=0$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\operatorname{dim\,Ker}\Omega$ $(U_{a},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=0$. And it is apparent that $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $\,(U_{a},H)$. Thus $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}$ $(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})$. According to Theorems 1 and 2, $(U_{a},H^{*}),\forall U_{a}\in\mathbb{R}^{k\times n}$ is an equilibrium whenever two players consider one-step or two-step optimality.
In this scenario, the defender can identify such an $H^{*}$ that makes $\mathcal{V}^{*}$ equal to zero. Then the attacker is unable to alter observability of the system any more and a trivial result is obtained. To avoid this special equilibrium case, set the game problem based on $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0.$ Then we want to give a condition for this setting.
Theorem 4.19.
For system (9), $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ and $H\in\mathbb{R}^{m\times n}$ are of full rank, if
$\max\{k,\text{MGM}({N}_{\bar{r}}))\}>m,$ (34)
where $\text{MGM}({N}_{\bar{r}})$ is the maximum geometric multiplicity associated with the eigenvalues of ${N}_{\bar{r}},$ ${N}_{\bar{r}}$ is the uncontrollable part for system matrix $N$, then $\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}>0.$
Proof 4.20.
Let the subscript $r$ represent the controllable part and the subscript $\bar{r}$ represent the uncontrollable part. The maximum controlled invariant subspace $\mathcal{V}^{*}$ can be divided into controllable part $\mathcal{V}^{*}_{r}$ and uncontrollable part $\mathcal{V}^{*}_{\bar{r}}$, where $\mathcal{V}^{*}_{r}$ is determined by $(N_{r},B^{\prime}_{ar},H_{r})$ and $\mathcal{V}^{*}_{\bar{r}}$ is determined by $(N_{\bar{r}},0,H_{\bar{r}})$.
(1) If $k>\text{MGM}({N}_{\bar{r}})$, then $k>m$. For controllable part $(N_{r},B^{\prime}_{ar},H_{r})$, the geometric multiplicity of $(N_{r}+B^{\prime}_{ar}U_{ar})^{\prime}s$ eigenvalue $\lambda$ is $\{n-\text{rank}[(N_{r}+B^{\prime}_{ar}U_{ar})-\lambda I]\}.$ According to PBH controllability criterion, for the controllable part, $\forall\lambda\in\sigma(N_{r}+B^{\prime}_{ar}U_{ar})$, rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)\,\,\,B^{\prime}_{ar}]=n$. Thus rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)]\geq(n-k)$ and $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})=k.$ If $k>m$, i.e., $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})>m$, according to Corollary 3.3, =$\mathcal{I}=\{i\mid\alpha_{i}>m\}\neq\emptyset,\underset{{H_{r}}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})=n-\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}})>0.$ According to Lemma 3, $\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}_{r}^{*}\geq\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})>0$. Thus $\min_{{H}_{r}}\operatorname{dim\,}$ $\mathcal{V}_{r}^{*}>0$.
(2) If $\text{MGM}({N}_{\overline{r}})>k$, then $\text{MGM}({N}_{\overline{r}})>m$. For uncontrollable subspace, $\mathcal{V}^{*}_{\bar{r}}=\operatorname{Ker}\Omega(N_{\bar{r}},0,H_{\bar{r}})$. If $\text{MGM}({N}_{\overline{r}})>m$, extending Lemma 2 to the dual system, we have $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,Ker}\Omega>0$. Thus $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0$.
In (1) or (2), $\operatorname{dim\,}\mathcal{V}^{*}=\operatorname{dim\,}\mathcal{V}_{r}^{*}+\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0.$
Under the conditions of $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0$, we obtain the following theorem.
Theorem 4.21.
For $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}$, if $\operatorname{dim\,}\mathcal{V}_{1}^{*}>0$ and $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*})$, then $\forall H\in\mathbb{R}^{m\times n},(U_{a}^{*},H)$ is not a Nash equilibrium.
Proof 4.22.
We prove it by the contrapositive: if $\exists H^{*}$ which makes $(U_{a}^{*},H^{*})$ a Nash equilibrium, then $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$, where $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\mathcal{V}^{*}(H)$. According to Lemma 3, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$. According to Theorems 1 and 2, if $(U_{a}^{*},H^{*})$ is a Nash equilibrium, $(U_{a}^{*},H^{*})$ satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,$ $(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$. Thus there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a}^{*},H)=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Therefore, we have $H^{*}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ and $\mathcal{V}^{*}(H^{*})=\mathcal{V}_{1}^{*}.$ Because $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,}\mathcal{V}_{1}^{*}$ according to Theorems 1 and 2, $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$. Thus the contrapositive has been proven.
4.4 Discussions
According to Theorems 1-3, we find that the defender dominates the game to achieve equilibrium. On the one hand, the defender can achieve equilibrium of the game by considering two-step optimality. This is because for fixed $U_{a}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H\in BR1^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR2^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR1^{d})$. $H\in BR2^{d}$ makes $\operatorname{dim\,}\mathcal{V}^{*}$ closer to $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ compared to $H\in BR1^{d}$. If they are equal, the condition for Nash equilibrium is satisfied. As for the attacker, it plays little role in generating equilibrium. Because for fixed $H_{1}$, $U_{a}^{*}\in BR1^{a}$ ((((or $U_{a}^{*}\in BR2^{a})$, there must be $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)\leq\operatorname{dim\,}\mathcal{V}^{*}(H_{1})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H_{1})$. Because the relation between $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)$ and $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})$ is unrelated to $U_{a}^{*}$, the attacker cannot make the game converge to equilibrium. On the other hand, Theorem 4.17 gives a case when the action of the defender and system matrices satisfy a specific condition, the game invariably remains in equilibrium regardless of the attacker’s action.
Considering Theorem 4.21, we find that the attacker can break the equilibrium regardless of the defender’s action if $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*}),$ where $\mathcal{V}_{1}^{*}$ is the smallest controlled invariant subspace. But in this case, the attacker does not maximize the current value function, forsaking the current gain. In fact, Nash equilibrium is sometimes disadvantageous to the attacker because the value function remains at a low value that the attacker cannot change. In this case, the attacker can break the equilibrium by choosing $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}^{*})$ and entice the defender to choose a lower value function action in the next epoch. However, the corresponding $\mathcal{V}^{*}$ value might be much higher, leading to the attacker achieving a much higher value function in the subsequent second epoch and in the future epochs. An example is given in Section 5 to illustrate this case.
To conclude, because the best response sets considering one-step optimality or two-step optimality are not single-valued maps, different choices in the best response sets will lead to different game results. Nevertheless, three rules of the game results can be summarized:
(1) The conditions for Nash equilibrium are the same whether players consider one-step optimality or two-step optimality.
(2) If both players are greedy, i.e., their actions ensure the optimum for the current epoch, the defender can find the Nash equilibrium following two-step optimality and has a dominant role in determining whether the game can reach an equilibrium.
(3) The attacker, by sacrificing current gains, can break the Nash equilibrium, enticing the defender to choose a different action associated with a lower value function and making the game become another pattern. Then the attacker may achieve a much higher value function in the future epochs.
In addition, this work only considers one-step optimality and two-step optimality, but does not consider three-step optimality or more steps, because it is the same player changing actions in the first step and in the third step. The game scenario in the first epoch is the same as that in the third epoch. So there is no need to consider higher steps.
5 Illustrative examples
In this section, we will illustrate the effectiveness of our main results using six cases. Consider a linear system
$\dot{z}$ $=Nz+B_{a}^{\prime}U_{a}z,$
$\hat{y}$ $=Hz,$
with $N={\left[\begin{array}[]{lllll}0.3&0&0&0&0\\ 0&0.3&0&0&0\\ 0&0&0.3&0&0\\ 0&0&0&0.1&0\\ 0&0&0&0&0.2\end{array}\right]},B_{a}^{\prime}=\left[\begin{array}[]{lllll}0&0&1&0&1\end{array}\right]^{\top}$. The attacker controls $U_{a}\in\mathbb{R}^{1\times 5}$ to maximize the unobservable subspace while the defender controls $H\in\mathbb{R}^{2\times 5}$ to minimize the unobservable subspace, forming an infinitely repeated game. In epoch 1, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$. The examples in this section are periodic. Although only the outcomes of 60 epochs are shown in the following figures, they are sufficient to represent the scenarios over an infinite time span.
Table 2: Game settings for Cases 1-6
Settings Attacker action rules Defender action rules Strategies
Case 1 Algorithm 1. Algorithm 2. One-step optimality
Case 2 $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]$ instead of $\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right],$ and other actions based on Algorithm 1. Algorithm 2.
Case 3 Algorithm 1. $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$ instead of $\!\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right],$ and other actions follow Algorithm 2.
Case 4 Before epoch 20: Algorithm 1; after 20 epochs: two-step optimality. Before epoch 40: Algorithm 2; after 40 epochs: two-step optimality. One-step & two-step optimality
Case 5 Before epoch 40: Algorithm 1; after 40 epochs: two-step optimality. Before epoch 20: Algorithm 2; after 20 epochs: two-step optimality.
Case 6 One-step optimality, only with $U_{a}\notin\mathcal{F}(\mathcal{V}^{*})$ in epoch 40. One-step optimality. One-step optimality & attacker not greedy
Table 3: Actions and average value functions of players for Cases 1-3.
Epochs $(n\in N^{*})$ Case 1 Case 2 Case 3
$n$ $H_{n}=\begin{bmatrix}1&0&0&1&1\\ 0&1&0&0&0\end{bmatrix}$ $H_{n}=\begin{bmatrix}1&0&0&1&1\\ 0&1&0&0&0\end{bmatrix}$ $H_{n}=\begin{bmatrix}1&0&0&1&1\\ 0&0&1&0&0\end{bmatrix}$
[…]
$\bar{\Phi}$ 1.5 2 1
The game settings for Case 1-6 are summarized in Table 2. Cases are divided into three categories (Cases 1-3; Cases 4-5; Case 6), which will be introduced separately below.
Cases 1-3: let players have different choices in the one-step optimal best response sets. In Case 1, the attacker chooses $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ according to Algorithm 1, which uses pseudo inverse to get $U_{a}$ with the minimum modulus length; the defender chooses $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$ according to Algorithm 2. While in Cases 2-3, players have different choices compared to Case 1.
Figure 5: Evolution of value function for Cases 1-3. (a) Case 1: the attacker uses Algorithm 1 and the defender uses Algorithm 2; (b) Case 2: the attacker changes strategy preference and other actions match Case 1; (c) Case 3: the defender changes strategy preference and other actions match Case 1.
Fig. 5 shows the evolution of unobservable subspace dimension for Cases 1-3, which are totally different in different cases. More specifically, the actions of players $U_{a}$ and $H$ evolve in different loops, which are summarized in Table 3. The average value functions $\bar{\Phi}$ are $1.5,2,1$ respectively for Cases 1-3.
The reasons for different utilities in different cases are as follows: in Case 1, $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]\in BR1^{a}$, but $U_{a}\notin BR2^{a};H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]\in BR1^{d}$, but $H\notin BR2^{d}$. In Case 2, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]\in BR2^{a}.$ In Case 3, $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]\in BR2^{d}$. As expected, when the action of the player falls exactly within its two-step optimal set, it can make the average value function more favorable for itself.
Cases 1-2 illustrate the results of Corollary 4.10 and 4.11. In Case 3, the value function and actions $U_{a}$ and $H$ are locked. Such $(U_{a},H)$ is a Nash equilibrium, which satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}(H)$. Case 3 illustrates the results of Theorem 4.5 and Corollary 4.8.
Figure 6: Game results for Cases 4-5. (a) Evolution of value function for Case 4; (b) Evolution of $\mathcal{V}^{*}$ space dimension for Case 4; (c) Evolution of value function for Case 5; (d) Evolution of $\mathcal{V}^{*}$ space dimension for Case 5.
Figure 7: Results of Case 6: game reaches equilibrium before epoch 40 and after epoch 40, the defender choose $U_{a}\notin BR1^{a}$. (a) Evolution of value function; (b) Evolution of $\mathcal{V}^{*}$ space dimension.
Cases 4-5: let two players start thinking about two-step optimality at different epochs to compare the impact of different players on the equilibrium. Before epoch 20, let two players consider one-step optimality and choose actions according to Algorithm 1 and Algorithm 2. In Case 4, let the attacker consider two-step optimality after 20 epochs and both players consider two-step optimality after 40 epochs. In Case 5, let the defender not the attacker consider two-step optimality after 20 epochs and other settings are the same as Case 4.
Fig. 6 shows the game results for Cases 4-5. Before epoch 20, the game results of both cases are the same as Case 1. From epoch 20-40, the value function evolves in a different loop in Case 4 and the game reaches the Nash equilibrium in Case 5. After epoch 40, the games in both cases reach the Nash equilibrium.
The reason why cases 4,5 reach equilibrium at different epochs is due to the dimension of $\mathcal{V}^{*}$. When the attacker considers two-step optimality at epoch 20, $\operatorname{dim}\mathcal{V}^{*}=3$, which is large and more beneficial to the attacker. In this case, $\operatorname{dim\,}\mathcal{V}^{*}>\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is no Nash equilibrium. When the defender considers two-step optimality, $\operatorname{dim}\mathcal{V}^{*}=1$, which is small and more beneficial to the defender. In this case $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is Nash equilibrium. And after epoch 40, there is $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$ in both cases, thus there is Nash equilibrium.
According to Cases 4-5, the defender can achieve Nash equilibrium by following two-step optimality but the attacker can not, and both cases illustrate the result of Theorem 4.15.
In the above five cases, both players are greedy, i.e., their actions are either one-step or two-step optimal. Next, consider a case when the attacker is not greedy.
Case 6: let both players consider one-step optimality and choose actions to achieve equilibrium before epoch 40. In epoch 40, when $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$, the attacker chooses $U_{a}=\left[\begin{array}[]{lllll}0&1&0&0&0\end{array}\right]\notin BR1^{a}$. Although the value function in epoch 40 reduces, $\operatorname{dim\,}\mathcal{V}^{*}$ in epoch 41 increases, which increases the value function in epoch 42. Thus the equilibrium is broken and the game reaches a new mode: the value function oscillates between 1 and 3. From the Nash equilibrium to the oscillation mode, the average value function increases from 1 to 1.5, which is more beneficial for the attacker. It shows that by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium and make the game result more beneficial for itself. This illustrates the result of Theorem 4.21.
6 Conclusion
In this paper, we formulate an infinitely repeated game model between the attacker and the defender based on the unobservable subspace dimension. Derivations and algorithms to maximize or minimize the unobservable subspace dimension are given. The game considering one-step optimality and two-step optimality are studied respectively. It is found that the best response sets are not single-valued maps and different choices will lead to different game results. Still, the necessary and sufficient condition for Nash equilibrium is given, which is the same no matter whether players consider one-step optimality or two-step optimality. We find if both players’ actions ensure the optimal of the current epoch, the defender has a dominant role in whether the game can reach an equilibrium. But by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium. Finally, six cases are given to illustrate the main results.
In the future, we can study the problem in more complex scenarios. Factors such as system stability, energy consumption and unobservable subspace can be combined to formulate a more comprehensive value function. Moreover, we can consider the game with incomplete information, for example with unknown system matrices, how does the attacker design strategies to change the observability of the system, and the strategy evolution of both sides.